OpenStudy (anonymous):

How many tangents are possible from the origin on the curve y=(x+1)^3 . Also find the equation of these tangents . HELP !

OpenStudy (anonymous):

this a calculus course?

OpenStudy (anonymous):

Yes , its a question from application of derivatives.

OpenStudy (anonymous):

a point on the curve looks like \((x, (x+1)^3)\) and the origin is \((0,0)\) the slope of the line between \((0,0)\) and \((x,(x+1)^3)\) is \(\frac{(x+1)^3}{x}\)

OpenStudy (anonymous):

the slope of the line tangent to the graph will be \(3(x+1)^2\)

OpenStudy (anonymous):

that being the derivative they have to be the same, set \[3(x+1)^2=\frac{(x+1)^3}{x}\] and solve for \(x\)

OpenStudy (anonymous):

Thanks a ton :)

OpenStudy (anonymous):

yw, hope it is clear more or less

OpenStudy (anonymous):

Yep , my answer matches :)